OpenStudy (anonymous):

Write 5.6 as a percent. 56% 5.6% 560% 0.56%

OpenStudy (jdoe0001):

hmm I'd think by now you would know, this about the 15th you've posted about %

OpenStudy (jdoe0001):

\(\bf {\color{red}{ a}}\%\ =\cfrac{{\color{red}{ a}}}{100}\) but this has been covered many times now

OpenStudy (anonymous):

b

OpenStudy (anonymous):

im not sure it is b and @jdoe0001 that did not work i only got 0.056 doing that

OpenStudy (anonymous):

Oh wait ik what it is now 560

OpenStudy (anonymous):

because u have to multiply it by 100

OpenStudy (anonymous):

5.6 is equal 10 5 6/10 right?

OpenStudy (jdoe0001):

\(\bf {\color{red}{ a}}\%=\cfrac{{\color{red}{ a}}}{100}\implies 100\times {\color{red}{ a}}\%={\color{red}{ a}}\\ \quad \\ {\color{red}{ 5.6}}\%=\cfrac{{\color{red}{ 5.6}}}{100}\implies 100\times {\color{red}{ 5.}}\%={\color{red}{ 5.6}}\)

OpenStudy (inkyvoyd):

@jdoe0001 , what??

OpenStudy (jdoe0001):

woops, I even missed a ... number ... dohh \(\bf {\color{red}{ a}}\%=\cfrac{{\color{red}{ a}}}{100}\implies 100\times {\color{red}{ a}}\%={\color{red}{ a}}\\ \quad \\ {\color{red}{ 5.6}}\%=\cfrac{{\color{red}{ 5.6}}}{100}\implies 100\times {\color{red}{ 5.6}}\%={\color{red}{ 5.6}}\)

OpenStudy (anonymous):

@jdoe0001 You are making it way harder then its supposed to be

OpenStudy (jdoe0001):

ohh hehehe... wait. a sec... I see what's amiss hheh

OpenStudy (jdoe0001):

alright... yes... indeed... one sec

OpenStudy (jdoe0001):

you'd just multiply the value given, as a plain number times 100 to get the % equivalent but you've done this many times now though

OpenStudy (jdoe0001):

\(\bf a=(a\times 100)\%\)

OpenStudy (anonymous):

look 5.6 equals 5 as a whole number and 6/10 equals five and six tenths. So therefore it is 560 because 6/10=60/100+500

OpenStudy (anonymous):

@MileenaGomez